OpenStudy (anonymous):

To produce 2.51 liters of a 1.55 M solution of magnesium nitrate, how many grams of magnesium nitrate must be dissolved?

OpenStudy (anonymous):

@aaronq

OpenStudy (aaronq):

there are a few ways to do this, i like substitution: \(n=\dfrac{m}{Molar~mass}\) into the molarity formula \(M=\dfrac{n_{solute}}{L_{solution}}\) \(M=\dfrac{\dfrac{m}{Molar~mass}}{L_{solution}} \rightarrow M=\dfrac{m}{L_{solution}*Molar~mass}\)

OpenStudy (aaronq):

plug in all the values (molarity, L, and Molar mass) then solve for mass, m.

OpenStudy (anonymous):

thanks:)

OpenStudy (aaronq):

no problem !